Roger's age is two times the age of his sister. last year his age was three times the age of his sister. How old is Roger?
Answer by mhel12jhane(6)

roger 2(x)-1
sister x-1
2x-1=3(x-1)
2x-1=3x-3
2x-3x=-3+1
-x=-2
x=2
roger is 4years old at present and 3 years old last year